OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. R(3,3) S(-6, -6) I got stuck on what to do next, heres what I have so far... -6 - 3 / -6 - 3 = -9/-9= 1 which is the slope y-3=1(x-3) y-3 = x -3 y=x +0

OpenStudy (amistre64):

well, youve got y=x correctly; standard form is: Ax + By = C ; such that A > 0

OpenStudy (amistre64):

well, A >= 0

OpenStudy (anonymous):

I asked my teacher and he said I have to move the x over so how do I do that?

OpenStudy (amistre64):

might want to consider moving the y over instead .... how do we get rid of something on one side and also "get rid of it" on the other side?

OpenStudy (amistre64):

consider what you did on this step: y-3 = x -3 ; you got rid of the 3 on one side, how did you get rid of it from the other side?

OpenStudy (amistre64):

y = x -y -y ; just subtract the same thing from both sides to keep things equal -------- 0 = x - y

OpenStudy (anonymous):

Oh my gosh thank you so much!

OpenStudy (amistre64):

it wasnt really a big deal :) youre welcome tho

OpenStudy (anonymous):

Well I was just stressing because I couldn't figure out what to do next!

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

Thanks!